# StarEmbed: Benchmarking Time Series Foundation Models on Astronomical Observations of Variable Stars

### v1

@@ 4.1 Unsupervised Clustering
We produce uncertainties on performance metrics by repeating the K-means algorithm with 10 different initializations, choosing the clustering with the lowest within-cluster variance. We also apply Ward’s hierarchical clustering, which optimizes the same within-cluster variance objective via agglomerative merges. It provides a robustness check since it is deterministic and initialization-free. We then evaluate clustering quality using the following literature-standard metrics for clustering (Z. Huang et al., 2020; T. Monnier et al., 2020; L. Sun et al., 2024; Y. Li et al., 2024):
@@
Macro-averaged F1 score (F1). Macro-F1 is the harmonic mean of the completeness (true positive rate) and purity (the false positive rate subtracted from unity) computed per class and then averaged evenly across all classes. Because it treats each class equally, the Macro-averaged F1 is sensitive to performance on minority classes. This is especially important for our CSPVS dataset because some classes (e.g., RRd) have many fewer labels than others (e.g., EW; see Table 1). Similarly to previous work (e.g., T. Monnier et al., 2020), we assign clusters to class predictions using the Hungarian matching algorithm (H. W. Kuhn, 1955). This treats clustering as unsupervised classification and makes results directly comparable to supervised settings.
@@ 4.2 Supervised Classification
To directly measure how useful the embeddings are for classifying variable stars, we train four simple classifiers on the fixed embeddings to predict the variable star class labels. This simulates a scenario where one uses a pretrained model to produce embeddings used as feature inputs for a classification task but does not fine-tune the embedding model (hence “zero-shot” in terms of the embedding model). We evaluate the embeddings with four (simple) classifiers: a non-parametric model ($k$-nearest neighbor, $k$-NN), a linear probe, a decision tree classifier (random forest, RF), and a non-linear model (multilayer perceptron, MLP). Both $k$-NN and linear probes are standard in the embedding evaluation literature (M. Caron et al., 2021; J. Zhou et al., 2021; A. Neelakantan et al., 2022), as they are simple methods that directly reflect separability in the embedding space. RF is included, in part, because it is widely used in the astronomical literature for periodic variable star classification (B. Naul et al., 2018; P. Sánchez-Sáez et al., 2021; Ó. Pimentel et al., 2022) and it achieves SOTA performance across many datasets (B. Naul et al., 2018). Finally, an MLP is used as a modern higher capacity deep-learning option. We report standard multi-class classification metrics to comprehensively assess the performance of each embedding model on the downstream supervised classification task. Accuracy: the fraction of stars that are correctly classified. Macro-averaged F1 Score (F1): The Macro-averaged F1 (see Sec. 4.1) is apt because it is sensitive to performance on minority classes. Precision/Recall: The overall precision and recall are included to provide a fully comprehensive evaluation of the classification performance.
@@
Hyperparameters. We report the hyperparameter tuning process and summary for all classifiers in this section. To fairly compare the different embeddings, we conduct a hyperparameter search on each model when training the downstream MLP and random forest classifier. We use MLP with three hidden layers of sizes 1024, 512 and 256 and an output layer for class predictions. we search over batch size $B\in\{128,256,512,1024\}$, learning rate $lr\in\{0.01,0.001,0.0001\}$ and dropout rate $\in\{0.0,0.1\}$. Every hyperparameter triple runs once on NVIDIA H100 4 GPUs. The training process is at most 50 epochs, and stops early if the validation loss fails to improve for 3 epochs. In practice this training takes less than 30 epochs for all models before the early stopping is triggered. For random forest, we search over maximum depth of the tree $\in\{\text{None},10,20,30\}$, the minimum number of samples to split an internal node $\in\{2,5,10\}$, and number of estimator $\in\{100,200,500\}$. We summarize the hyperparameters of the MLP and random forest classifiers in Table 2 and  3. For linear classifier, since the current training set is relatively small, we use the LogisticRegression (L-BFGS) from Scikit-learn library (L. Buitinck et al., 2013), with max_iter = 5000, class_weight = "balanced", and all other with default settings. L-BFGS is a deterministic full-batch method that converges to the global minimizer without learning rate tuning. For $k$-NN, we use KNeighborsClassifier from Scikit-learn with default settings. Since $k$-NN and the default solver of logistic regression are both deterministic, we only report the 1 run result.
4.3 Out-of-distribution source detection
@@
Identifying variable stars physically unlike those in labeled training sets is of great astrophysical-interest. To test the effectiveness of the embeddings for detecting such OOD sources, we compute OOD scores for light curves with a modified isolation forest algorithm (R. Gupta et al., 2025). We first create embeddings for the ZTF light curves of CSPVS stars with too few examples to be included in the training set ($\beta$-Lyrae, Blazhko, Anomalous Cepheids, Cepheid-II, HADS, LADS, ELL, Hump, PCEB, and EAup; see Sec. 3). We define these as OOD sources. The embeddings of the OOD sources are mixed with the test set and run through a “multi-class isolation forest" (R. Gupta et al., 2025) where a separate isolation forest is fit to the embeddings of each of the seven inlier classes in the training set. The minimum of the seven isolation forest scores is the OOD score we use for OOD source detection. Isolation forest is a popular method for finding astrophysical outliers (K. L. Malanchev et al., 2021b), and R. Gupta et al. (2025) show that following this multi-class prescription yields better macro-averaged performance than a single isolation forest in many settings, including for periodic variables. Here, the performance is benchmarked with the fraction of sources in the top $N^{\textrm{th}}$ percentile of OOD scores which are genuine OOD sources: the OOD purity.
5 Benchmark Results
@@ 5.1 Unsupervised Clustering
Astromer-1 0.0041(0.0002) 0.0022(0.0011) 0.1659(0.0022) 0.0041 0.0001 0.1652
Astromer-2 0.0076(0.0003) 0.0143(0.0002) 0.1571(0.0030) 0.0091 0.0310 0.1600
Moirai-small 0.1724(0.0010) 0.0978(0.0059) 0.2816(0.0068) 0.1476 0.0828 0.2612
Chronos-tiny 0.2432(0.0000) 0.1590(0.0001) 0.3410(0.0000) 0.1890 0.1217 0.3671
Chronos-Bolt-tiny 0.2134(0.0004) 0.1245(0.0011) 0.3136(0.0001) 0.2273 0.1553 0.3662
Random Embeddings 0.0003(0.0001) 0.0000(0.0000) 0.0971(0.0014) 0.0003 0.0004 0.1122
Hand-crafted Features 0.2619(0.0058) 0.1139(0.0071) 0.4028(0.0277) 0.2508 0.1319 0.3323
@@
5.2 Supervised Classification
Table 6 and the left panel of Figure 2 show that (i) the Chronos models once again perform very well compared to Moirai-small and the Astromer models; (ii) Astromer-2 performs better than Moirai-small in some metrics; and (iii) unlike in the clustering results, Chronos-tiny outperforms Chronos-Bolt-tiny and achieves the second best performance in nearly all metrics. As in the clustering results, the hand-crafted features are clearly superior, yielding a F1 score of $0.807\pm 0.002$ with the RF classifier. We also find that the RF classifier generally performs better than others, although this is somewhat model-dependent.
The center and right panels of Figure 2 show the confusion matrix of one of the best performing TSFM-classifier pairings (Chronos-tiny with the MLP) and best overall performing model-classifier pairing (hand-crafted features with an RF). These confusion matrices show that (i) both Chronos-tiny and the hand-crafted features often confuse RRd sources as RRc; (ii) Chronos-tiny yields better performance on most classes (EA, RRd, RS CVn, and LPV) although loses overall due to the larger margins in the classes where the hand-crafted features perform better (EW, RRab, RRc). In general, these results show that, despite having never seen astronomical time series, Chronos-tiny clearly extracts useful information from the data for supervised classification. The complete set of confusion matrices across all embedding–classifier combinations is presented in Appendix D.
@@ 5.2 Supervised Classification
5.3 Out-of-distribution source detection
@@
Astromer-1 0.014 (0.020) 0.087 (0.018) 0.121 (0.001)
Astromer-2 0.139 (0.019) 0.124 (0.001) 0.119 (0.001)
Moirai-small 0.172 (0.023) 0.143 (0.003) 0.152 (0.002)
Chronos-tiny 0.165 (0.005) 0.129 (0.017) 0.158 (0.021)
Chronos-Bolt-tiny 0.569 (0.037) 0.536 (0.053) 0.528 (0.050)
@@
Hand-crafted features 0.213(0.009) 0.280(0.006) 0.260(0.002)
@@
RF Accuracy 0.676 (0.000) 0.846 (0.000) 0.822 (0.001) 0.863 (0.000) 0.827 (0.000) 0.676 (0.000) 0.921 (0.000)
Precision 0.144 (0.067) 0.794 (0.005) 0.706 (0.002) 0.771 (0.070) 0.705 (0.003) 0.097 (0.000) 0.865 (0.001)
Recall 0.143 (0.000) 0.525 (0.003) 0.512 (0.002) 0.599 (0.000) 0.548 (0.002) 0.143 (0.000) 0.777 (0.002)
F1 0.115 (0.000) 0.580 (0.003) 0.554 (0.001) 0.640 (0.001) 0.582 (0.002) 0.115 (0.000) 0.807 (0.002)
MLP Accuracy 0.459 (0.162) 0.535 (0.106) 0.711 (0.045) 0.775 (0.017) 0.740 (0.013) 0.338 (0.157) 0.842 (0.011)
Precision 0.143 (0.008) 0.372 (0.036) 0.565 (0.019) 0.594 (0.001) 0.569 (0.010) 0.137 (0.024) 0.684 (0.030)
Recall 0.162 (0.003) 0.538 (0.041) 0.694 (0.014) 0.760 (0.006) 0.716 (0.004) 0.146 (0.002) 0.859 (0.001)
F1 0.134 (0.016) 0.388 (0.051) 0.592 (0.013) 0.645 (0.008) 0.612 (0.009) 0.096 (0.047) 0.735 (0.023)
5.4 Visualizations of Embeddings
We include the UMAP visualizations for the embeddings from each embedding models to provide more intuitions regarding the embedding space. As shown by Figure 3, all time series pretrained models’ embeddings are showing clear distinction and distribution of different clusters corresponding to different ground truth classes. In comparison, as a baseline, the random embeddings show no clear clusters at all. Astromer-1’s embeddings and hand crafted features do not show clear clusters either. Astromer-2’s embeddings show clearer cluster distribution but for some classes, the clusters are not distinctive with others either. These UMAP visualizations further demonstrate the promising potentials of using time series pretrained models as light curve embedding models.
@@ 5.4 Visualizations of Embeddings
6 Discussion and Conclusions
We have introduced StarEmbed, a public benchmark for evaluating time series foundation models (TSFMs) on observed multi-band stellar light curves, which pose unique challenges, irregular sampling and heteroskedasticity, relative to typical TSFM pretraining corpora. By harmonizing expert-vetted CSPVS labels with $\sim$40,000 multi-band ZTF light curves, we deliver a rigorously curated, seven-class dataset together with a thorough evaluation of three core tasks in light curve analysis: unsupervised clustering, variable-star classification, and OOD detection. Using this benchmark, we perform a comprehensive comparison of (i) domain-specific embeddings (the Astromer models), (ii) embeddings from three SOTA general-purpose TSFMs (Moirai and Chronos families), and (iii) hand-crafted features as the primary baseline. Three main findings emerge: (i) for clustering, Chronos provides comparable performance to hand-crafted features; (ii) for supervised classification, hand-crafted features outperform all other models, while Chronos is consistently second, trailing the hand-crafted features by a small margin; and (iii) for OOD detection Chronos-Bolt-tiny significantly outperforms hand-crafted features. For all three tasks the TSFMs almost always outperform the Astromer models in zero-shot.
Taken together, our results demonstrate the limits of TSFMs’ generalization ability and advocate for a shift from bespoke, fully-supervised pipelines toward off-the-shelf foundation representations plus lightweight task heads for variable-star analysis and, by extension, for other forthcoming petascale time-series challenges. By releasing all data, code, and model wrappers, we hope StarEmbed to serve as a community reference benchmark for self-supervised and foundation-model advances, spur work on even broader time-series corpora, and catalyze progress on downstream tasks such as period estimation and real-time alert triage in next-generation sky surveys like LSST.
7 Acknowledgement

### v2

@@ IV.1 Unsupervised Clustering
We produce uncertainties on performance metrics by repeating the K-means algorithm with 10 different initializations, choosing the clustering with the lowest within-cluster variance. We also apply Ward’s hierarchical clustering, which optimizes the same within-cluster variance objective via agglomerative merges. It provides a robustness check since it is deterministic and initialization-free. We then evaluate clustering quality using the following literature-standard metrics for clustering (Huang et al., 2020; Monnier et al., 2020; Sun et al., 2024; Li et al., 2024):
@@
Macro-averaged F1 score (F1). Macro-F1 is the harmonic mean of the completeness (true positive rate) and purity (the false positive rate subtracted from unity) computed per class and then averaged evenly across all classes. Because it treats each class equally, the Macro-averaged F1 is sensitive to performance on minority classes. This is especially important for our CSPVS dataset because some classes (e.g., RRd) have many fewer labels than others (e.g., EW; see Table 1). Similarly to previous work (e.g., Monnier et al., 2020), we assign clusters to class predictions using the Hungarian matching algorithm (Kuhn, 1955). This treats clustering as unsupervised classification and makes results directly comparable to supervised settings.
@@ IV.2 Supervised Classification
To directly measure how useful the embeddings are for classifying variable stars, we train four simple classifiers on the fixed embeddings to predict the variable star class labels. This simulates a scenario where one uses a pretrained model to produce embeddings used as feature inputs for a classification task but does not fine-tune the embedding model (hence “zero-shot” in terms of the embedding model). We evaluate the embeddings with four (simple) classifiers: a non-parametric model ($k$-nearest neighbor, $k$-NN), a linear probe, a decision tree classifier (random forest, RF), and a non-linear model (multilayer perceptron, MLP). Both $k$-NN and linear probes are standard in the embedding evaluation literature (Caron et al., 2021; Zhou et al., 2021b; Neelakantan et al., 2022), as they are simple methods that directly reflect separability in the embedding space. RF is included, in part, because it is widely used in the astronomical literature for periodic variable star classification (Naul et al., 2018; Sánchez-Sáez et al., 2021; Pimentel et al., 2022) and it achieves SOTA performance across many datasets (Naul et al., 2018). Finally, an MLP is used as a modern higher capacity deep-learning option. We report standard multi-class classification metrics to comprehensively assess the performance of each embedding model on the downstream supervised classification task. Accuracy: the fraction of stars that are correctly classified. Macro-averaged F1 Score (F1): The Macro-averaged F1 (see Sec. IV.1) is apt because it is sensitive to performance on minority classes. Precision/Recall: The overall precision and recall are included to provide a fully comprehensive evaluation of the classification performance.
@@
Hyperparameters. We report the hyperparameter tuning process and summary for all classifiers in this section. To fairly compare the different embeddings, we conduct a hyperparameter search on each model when training the downstream MLP and random forest classifier. We use MLP with three hidden layers of sizes 1024, 512 and 256 and an output layer for class predictions. we search over batch size $B\in\{128,256,512,1024\}$, learning rate $lr\in\{0.01,0.001,0.0001\}$ and dropout rate $\in\{0.0,0.1\}$. Every hyperparameter triple runs once on NVIDIA H100 4 GPUs. The training process is at most 50 epochs, and stops early if the validation loss fails to improve for 3 epochs. In practice this training takes less than 30 epochs for all models before the early stopping is triggered. For random forest, we search over maximum depth of the tree $\in\{\text{None},10,20,30\}$, the minimum number of samples to split an internal node $\in\{2,5,10\}$, and number of estimator $\in\{100,200,500\}$. We summarize the hyperparameters of the MLP and random forest classifiers in Table 2 and  3. For linear classifier, since the current training set is relatively small, we use the LogisticRegression (L-BFGS) from Scikit-learn library (Buitinck et al., 2013), with max_iter = 5000, class_weight = "balanced", and all other with default settings. L-BFGS is a deterministic full-batch method that converges to the global minimizer without learning rate tuning. For $k$-NN, we use KNeighborsClassifier from Scikit-learn with default settings. Since $k$-NN and the default solver of logistic regression are both deterministic, we only report the 1 run result.
IV.3 Out-of-distribution source detection
@@
Identifying variable stars physically unlike those in labeled training sets is of great astrophysical-interest. To test the effectiveness of the embeddings for detecting such OOD sources, we compute OOD scores for light curves with a modified isolation forest algorithm (Gupta et al., 2025). We first create embeddings for the ZTF light curves of CSPVS stars with too few examples to be included in the training set ($\beta$-Lyrae, Blazhko, Anomalous Cepheids, Cepheid-II, HADS, LADS, ELL, Hump, PCEB, and EAup; see Sec. III). We define these as OOD sources. The embeddings of the OOD sources are mixed with the test set and run through a “multi-class isolation forest" (Gupta et al., 2025) where a separate isolation forest is fit to the embeddings of each of the seven inlier classes in the training set. The minimum of the seven isolation forest scores is the OOD score we use for OOD source detection. Isolation forest is a popular method for finding astrophysical outliers (Malanchev et al., 2021b), and Gupta et al. (2025) show that following this multi-class prescription yields better macro-averaged performance than a single isolation forest in many settings, including for periodic variables. Here, the performance is benchmarked with the fraction of sources in the top $N^{\textrm{th}}$ percentile of OOD scores which are genuine OOD sources: the OOD purity.
V Benchmark Results
@@ V.1 Unsupervised Clustering
Astromer-1 0.0041(0.0001) 0.0017(0.0011) 0.1660(0.0014) 0.0041 0.0001 0.1652
Astromer-2 0.0082(0.0010) 0.0192(0.0078) 0.1590(0.0042) 0.0091 0.0310 0.1600
Moirai-small 0.1749(0.0017) 0.0981(0.0028) 0.2831(0.0034) 0.1476 0.0828 0.2612
Chronos-tiny 0.2374(0.0082) 0.1596(0.0029) 0.3110(0.0362) 0.1890 0.1217 0.3671
Chronos-Bolt-tiny 0.2120(0.0033) 0.1306(0.0125) 0.3128(0.0027) 0.2273 0.1553 0.3662
Random Embeddings 0.0003(0.0001) 0.0000(0.0000) 0.0977(0.0007) 0.0003 0.0004 0.1122
Hand-crafted Features 0.2700(0.0058) 0.1197(0.0092) 0.3960(0.0271) 0.2508 0.1319 0.3323
@@
V.2 Supervised Classification
Table 6 and the left panel of Figure 3 show that (i) the Chronos models once again perform very well compared to Moirai-small and the Astromer models; (ii) Astromer-2 performs better than Moirai-small in some metrics; and (iii) unlike in the clustering results, Chronos-tiny outperforms Chronos-Bolt-tiny and achieves the second best performance in nearly all metrics. As in the clustering results, the hand-crafted features are clearly superior, yielding a F1 score of $0.804\pm 0.003$ with the RF classifier. We also find that the RF classifier generally performs better than others, although this is somewhat model-dependent. We conduct a feature importance analysis on the hand-crafted features in Appendix D and find that the period of the variability in either passband are the most important features.
The center and right panels of Figure 3 show the confusion matrix of one of the best performing TSFM-classifier pairings (Chronos-tiny with the MLP) and best overall performing model-classifier pairing (hand-crafted features with an RF). These confusion matrices show that (i) both Chronos-tiny and the hand-crafted features often confuse RRd sources as RRc; (ii) Chronos-tiny yields better performance on most classes (EA, RRd, RS CVn, and LPV) although loses overall due to the larger margins in the classes where the hand-crafted features perform better (EW, RRab, RRc). In general, these results show that, despite having never seen astronomical time series, Chronos-tiny clearly extracts useful information from the data for supervised classification. The complete set of confusion matrices across all embedding–classifier combinations is presented in Appendix G.
@@ V.2 Supervised Classification
V.3 Out-of-distribution source detection
@@
Astromer-1 0.017(0.016) 0.091(0.019) 0.120(0.001)
Astromer-2 0.135(0.027) 0.126(0.006) 0.120(0.002)
Moirai-small 0.169(0.024) 0.143(0.007) 0.150(0.004)
Chronos-tiny 0.139(0.054) 0.116(0.028) 0.149(0.021)
Chronos-Bolt-tiny 0.569(0.060) 0.532(0.055) 0.519(0.038)
@@
Hand-crafted features 0.213(0.013) 0.271(0.014) 0.259(0.003)
@@
RF Accuracy 0.676 (0.000) 0.846 (0.000) 0.823 (0.001) 0.862 (0.000) 0.826 (0.001) 0.676 (0.000) 0.920 (0.001)
Precision 0.111 (0.043) 0.799 (0.006) 0.716 (0.007) 0.750 (0.056) 0.707 (0.002) 0.097 (0.000) 0.866 (0.003)
Recall 0.143 (0.000) 0.526 (0.002) 0.514 (0.002) 0.597 (0.002) 0.548 (0.001) 0.143 (0.000) 0.773 (0.004)
F1 0.115 (0.000) 0.580 (0.002) 0.557 (0.002) 0.638 (0.002) 0.582 (0.001) 0.115 (0.000) 0.804 (0.003)
MLP Accuracy 0.446 (0.147) 0.627 (0.037) 0.717 (0.031) 0.783 (0.022) 0.721 (0.022) 0.308 (0.203) 0.833 (0.022)
Precision 0.154 (0.006) 0.453 (0.019) 0.546 (0.022) 0.589 (0.025) 0.553 (0.026) 0.137 (0.006) 0.672 (0.025)
Recall 0.165 (0.003) 0.627 (0.020) 0.722 (0.006) 0.758 0.006) 0.696 (0.013) 0.145 (0.002) 0.851 (0.009)
F1 0.138 (0.020) 0.470 (0.023) 0.594 (0.019) 0.643 (0.023) 0.589 (0.015) 0.094 (0.044) 0.723 (0.027)
V.4 Visualizations of Embeddings
We include UMAP visualizations for embeddings from each embedding models to provide more intuitions regarding the embedding space. As shown by Figure 4, all time series pretrained models’ embeddings are showing clear distinction and distribution of different clusters corresponding to different ground truth classes. In comparison, as a baseline, the random embeddings show no clear clusters at all. Astromer-1’s embeddings and hand crafted features do not show clear clusters either. Astromer-2’s embeddings show clearer cluster distribution but for some classes, the clusters are not distinctive with others either. These UMAP visualizations further demonstrate the promising potentials of using time series pretrained models as light curve embedding models.
@@ V.4 Visualizations of Embeddings
VI Discussion and Conclusions
Figure 5: A long period variable (LPV) star’s light curve indexed by the measurement timestamps (left: the real, physical index) and indexed by the order of the measurements (right: a non-physical index). In both panels, the coloring corresponds to the physical time index, showing that discarding relative timestamp information causes the TSFMs to receive a warped, view of the light curve, limiting their interpretation of the data.
We introduced StarEmbed, a public benchmark for evaluating TSFMs on real multi-band stellar light curves, whose irregular sampling and heteroskedasticity differ substantially from typical TSFM pretraining corpora. By harmonizing expert-vetted CSPVS labels with $\sim$40,000 ZTF light curves, we provide a rigorously curated seven-class dataset and evaluate three key tasks: unsupervised clustering, variable-star classification, and OOD detection. On this benchmark, we compare (i) domain-specific embeddings (Astromer); (ii) SOTA general-purpose TSFMs (Moirai and Chronos); and (iii) hand-crafted features. Three findings emerge: (i) for clustering, Chronos matches hand-crafted features; (ii) for supervised classification, hand-crafted features perform best, with Chronos consistently second by a small margin; and (iii) for OOD detection, Chronos-Bolt-tiny significantly outperforms hand-crafted features. Across all tasks, TSFMs generally outperform Astromer in zero-shot.
Taken together, our results demonstrate the boundaries of TSFM generalization. Figure 5 illustrates a key limitation of TSFMs: not properly treating irregularly sampled data results in a significantly warped view of the time series. This warped view destroys key information like the period of the variability which we found to be the most important feature driving the performance of our strong baseline (Appendix D). TSFMs which (i) introduce architectural mechanisms to treat irregularly sampled time series and (ii) include such data in their pretraining corpora will alleviate this problem and likely excel in our benchmark. In astronomy, such a model would drive a paradigm shift from bespoke, fully-supervised pipelines toward off-the-shelf foundational representations plus lightweight task heads for variable star analysis. These frameworks would push analysis frontiers in the, now ongoing, era of astronomy driven by observatories creating petabyte-scale datasets (Ivezić et al., 2019). In response, astronomy has challenging, information-rich time series in game-changing quantities to offer the AI for time series field. Making use of these time series would also help open new applications of TSFMs beyond astronomy as other scientific domains also produce irregularly sampled time series. By releasing all data, code, and model wrappers, StarEmbed serves to forge the bridge between astronomy and the AI for time series fields and to provide the community with a benchmark for foundation model advances on challenging time series data.
VII Acknowledgement